OpenStudy (anonymous):

HELP!! suppose that x=x(t) and y=y(t) are both functions of t. if x^2+y^2=10, and dx/dt = -4 when x=3 and y =1, what is dy/dt?

OpenStudy (dumbcow):

using implicit differentiation you get \[2x \frac{dx}{dt} + 2y \frac{dy}{dt} = 0\] solving for dy/dt \[\frac{dy}{dt} = -\frac{x \frac{dx}{dt}}{y}\]

OpenStudy (anonymous):

so do i just plug in now?

OpenStudy (anonymous):

i got -12/1 but it isnt correct

OpenStudy (anonymous):

nvm